OpenStudy (anonymous):

Identify two arithmetic means between -31 and -10.

OpenStudy (nikita2):

-31 + -10 /2 = -41/2 = -20.5 is it?

OpenStudy (anonymous):

i dont see how there could be 2 arithmetic means >.< the arithmetic mean of two numbers (say a and b) is: \[\frac{a+b}{2}\] so the answer nikita posted is correct.

OpenStudy (anonymous):

-31 , x , y, -10 2x=-31+ y ....(i) 2y = -10 + x ...........(ii) (ii) solve for x x = 2y + 10 Substitute for x in (i) 2(2y + 10) = -31+ y 4y + 20 = -31 + y 3y = -51 y = -17 x = 2(-17) + 10 = -34+ 10 = -24 the 2 means are -24 and -17

OpenStudy (anonymous):

Thank you very much!